OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y + 3 = -6(x + 2) y + 2 = 6(x + 3) y - 3 = -6(x - 2) y - 2 = 6(x - 3)

OpenStudy (anonymous):

@Leader

OpenStudy (anonymous):

@pooja195

OpenStudy (leader):

Hi there :)

Nnesha (nnesha):

i guess u asked this question yesterday right

OpenStudy (anonymous):

is this a?

OpenStudy (anonymous):

no @Nnesha

OpenStudy (leader):

And since you know that the slope is -6, there are only 2 options iwth that. A or C.

Nnesha (nnesha):

yes good job

OpenStudy (leader):

One good way to figure it out is once you have found the slope, plug in the 2 points and see if it works with the equation :)

OpenStudy (anonymous):

The graph of a function f(x) is shown below: graph of line going through ordered pairs -2, -1 and 3, 3. An open circle at -2, -1 and a black dot at 3, 3 What is the domain of f(x)? -2 < x ≤ 3 -2 ≤ x < 3 -1 < y ≤ 3 -1 ≤ y < 3

OpenStudy (anonymous):

c?

OpenStudy (leader):

@heymydearpeople My bad, for the first question it is A. :) Good Job :P

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@Leader am i right ?

Nnesha (nnesha):

told u u asked this question yesterday :P :D

OpenStudy (leader):

It can not be C. It asks for domain and you have y in your answer. It has to be x.

OpenStudy (anonymous):

A sequence has its first term equal to 4, and each term of the sequence is obtained by adding 2 to the previous term. If f(n) represents the nth term of the sequence, which of the following recursive functions best defines this sequence? f(1) = 2 and f(n) = f(n - 1) + 4; n > 1 f(1) = 4 and f(n) = f(n - 1) + 2n; n > 1 f(1) = 2 and f(n) = f(n - 1) + 4n; n > 1 f(1) = 4 and f(n) = f(n - 1) + 2; n > 1

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

@Nnesha @sleepyjess @Leader, i ahem to leave soon :(

OpenStudy (leader):

Like @sleepyjess said in the other question. its not +2n, its +2 :)